### v1

# SAM: A Stability-Aware Cache Manager for Multi-Tenant Embedded Databases

Abstract.
The co-location of multiple database instances on resource constrained edge nodes creates significant cache contention, where traditional schemes are inefficient and unstable under dynamic workloads. To address this, we present SAM, an autonomic cache manager powered by our novel AURA algorithm. AURA makes stability a first-class design principle by resolving the exploitation-exploration dilemma: it achieves this by synthesizing two orthogonal factors, which we introduce as: the $\mathcal{H}$-factor, representing a database’s proven, historically stable efficiency (exploitation), and the $\mathcal{V}$-factor, representing its empirically estimated marginal gain for future improvements (exploration). This dual-factor model, governed by an adaptive weight, enables SAM to achieve sustained high performance through strategic stability and robustness in volatile conditions.
Extensive experiments against 14 diverse baselines demonstrate SAM’s superiority. It achieves top-tier throughput while being uniquely resilient to complex workload shifts and cache pollution attacks. Furthermore, its decision latency is highly scalable, remaining nearly constant as the system grows to 120 databases. Crucially, SAM achieves superior decision stability—maintaining consistent optimization directions despite noise, avoiding performance oscillations while ensuring predictable Quality of Service. These results prove that a principled, stability-aware design is essential for sustained high performance in real-world, large-scale systems.
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://github.com/maxsignalll/SAM.
1. Introduction
Embedded databases, such as SQLite (Hipp, 2024b), are integral to the modern computing ecosystem. Unlike traditional client-server (C/S) databases (Group, 2024; Corporation, 2024), they are designed as libraries to be directly integrated into an application. This architecture excels in scenarios requiring local, high-performance, and zero-management data storage, including mobile applications (Developers, 2024a), IoT devices (Project, 2024), and metadata management in large-scale distributed systems (Borthakur et al., 2013). Statistically, embedded databases are the most widely deployed database engines (Hipp, 2024a), rendering it nearly ubiquitous.
However, the increasing complexity of systems has given rise to a critical challenge: the concurrent operation of multiple, independent, database-reliant services on a single host, a practice known as multi-tenant co-location, where each service and its database acts as a distinct tenant. An edge computing node, for example, may simultaneously run services for network monitoring, data acquisition, and local analytics, each employing its own SQLite database. Similarly, in cloud-native environments, a physical machine may host dozens of containerized tenants that use SQLite for state management, a pattern increasingly advocated by modern cloud platforms for deploying scalable, distributed services (Developers, 2024b; Documentation, 2024).
In such multi-tenant co-location environments, a significant performance bottleneck emerges from memory contention—specifically, for the cache of each database instance. The conventional approach of statically assigning a fixed cache size to each instance is profoundly inefficient under dynamic workloads (Mehta and DeWitt, 1993). An idle instance’s cache lies fallow, wasting memory, while an instance experiencing a sudden load spike is starved for resources, leading to unpredictable performance swings that prevent sustained high throughput.
This inefficiency is particularly critical in modern Cloud-Edge-Device architectures. Here, cache misses trigger expensive network requests to remote data centers rather than local disk accesses, resulting in latencies that are orders of magnitude higher. Consequently, poor cache management creates volatile performance patterns, making sustained high performance impossible and user experience unpredictable.
Addressing this bottleneck is complicated by the content-dependent and non-linear nature of cache performance (Zhang et al., 2020; Berger et al., 2018), which creates a complex optimization problem with inherently conflicting objectives:
  • (C1) Workload Dynamism: The system must be agile enough to swiftly reallocate cache to new, emerging hotspots as workloads shift (Chen et al., 2020).
  • (C2) Cache Pollution: It must be discerning enough to prevent “cache polluters”—instances running cache-unfriendly workloads like full table scans that consume vast amounts of memory for little gain—from evicting more valuable data (Seshadri et al., 2012; Mittal, 2017).
  • (C3) SLA Guarantees: It must ensure that low-traffic but high-priority instances are not starved of resources by busier, less-critical tenants (Berger et al., 2018).
  • (C4) Decision Stability: Finally, it must maintain consistent optimization decisions despite transient noise (e.g., network jitter) while remaining responsive to genuine workload shifts, as performance oscillations severely degrade sustained throughput (Yang et al., 2022).
These requirements often pull in opposite directions—agility conflicts with stability, protection conflicts with efficiency. Addressing these conflicting requirements in unison necessitates a strategy that considers multiple orthogonal factors. As a structural foundation, we adapt the mature “reservation plus elasticity” model from cloud computing, which we instantiate in our system as a two-pool architecture (Liakopoulos et al., 2019; Al-Dhuraibi et al., 2017). This design provides each tenant with a guaranteed resource baseline via a dedicated Fixed Pool (the reservation) while enabling dynamic adjustments from a shared Elastic Pool (the elasticity). This architecturally provides a direct solution for (C3) SLA Guarantees and creates the foundation for addressing the remaining challenges through principled control of the elastic pool.
However, the model itself is incomplete, as a naive policy for the elastic pool fails to resolve the remaining dynamic challenges. A purely reactive policy may struggle with (C1) Workload Dynamism, while a purely opportunistic one is vulnerable to (C2) Cache Pollution and can induce severe performance oscillations (C4). Therefore, the core problem this paper addresses is the design of an adaptive control policy for this elastic pool that achieves sustained high performance in volatile environments. Our key insight is that such performance requires decision stability as its foundation, which can only be achieved by explicitly navigating the fundamental trade-off between rewarding proven, historical efficiency (exploitation) and seeking future performance improvements (exploration).
To effectively balance this exploitation-exploration trade-off, we propose SAM (a Stability-Aware Manager), an autonomic cache management system powered by our novel AURA (Autonomic Utility-balancing Resource Allocator) algorithm. As illustrated in Figure 1, SAM’s design directly maps onto the challenges. First, its two-pool allocation model institutionalizes the “reservation plus elasticity” paradigm, providing deterministic baseline resources via a fixed pool to directly address (C3) SLA Guarantees. Second, and most critically, the AURA algorithm provides the principled, dual-factor scoring model for the elastic pool. It balances a database’s historically stable efficiency ($\mathcal{H}$-factor) to prevent (C2) Cache Pollution, with its estimated marginal gain ($\mathcal{V}$-factor) to adapt to (C1) Workload Dynamism. Finally, AURA’s stability-enforcing actuation layer ensures (C4) Decision Stability. Together, these mechanisms enable SAM to achieve sustained high performance by combining decision stability with robustness against external disturbances.
The main contributions of this paper are threefold as follows.
  1. (1)
    We propose SAM, a novel cache management system for multi-tenant embedded databases. Its two-pool architecture is the first to systematically decouple baseline SLA guarantees from dynamic performance optimization.
  2. (2)
    We propose the AURA algorithm, whose novelty lies in a dual-factor model that balances a tenant’s historically stable efficiency ($\mathcal{H}$-factor) with its forward-looking marginal gain ($\mathcal{V}$-factor). This principled synthesis achieves decision stability that prevents self-inflicted performance oscillations and robustness against deceptive workloads, together enabling sustained high performance in volatile environments.
  3. (3)
    Through extensive evaluation against 14 diverse baselines, we demonstrate that SAM achieves top-tier throughput while maintaining unique robustness to adversarial workloads like cache pollution. Our scalability analysis shows near-constant decision latency as the system scales to 120 databases.
The remainder of this paper is organized as follows. Section 2 presents the system model and problem formulation. Section 3 details the SAM system’s architecture and design philosophy. Section 4 presents the AURA scheduling algorithm in detail. Section 5 provides a theoretical analysis of the algorithm’s stability properties and its computational complexity. Section 6 describes our experimental setup and evaluation. Section 7 reviews related work, and Section 8 concludes.
2. Problem Definition
In this section, we formalize the problem of cache management for multi-tenant embedded databases. We define the system architecture, the performance metrics of interest, and the multi-objective optimization problem that our framework, SAM, is designed to solve.
2.1. System Model
Our system model comprises a set of $N$ independent database instances, $D=\{d_{1},...,d_{N}\}$, co-located on a single host. These instances compete for a shared, finite total cache budget, $C_{total}$, which is measured in pages. The core problem is to dynamically determine the cache allocation for each instance, $a_{i}(t)$, at each time interval $t$, such that $\sum_{i=1}^{N}a_{i}(t)=C_{total}$. The performance of each instance is a function of its allocation, and we monitor key metrics such as its workload intensity (Operations per Second, opsi) and its cache effectiveness (Hit Rate, $HR_{i}$).
2.2. Problem Formulation
The cache allocation problem in our multi-tenant embedded databases environment is fundamentally a multi-objective optimization problem: an ideal allocation strategy must maximize system throughput while ensuring tail latency remains below specified SLA thresholds. This multi-objective goal can be formalized as maximizing a utility function $U(\mathbf{TPS},\mathbf{L_{p99}})$, but directly optimizing such a complex function in a noisy, online setting is intractable.
We therefore posit that Aggregate Effective Throughput is a powerful and stable proxy metric for this overarching goal. In the Cloud-Edge-Device architectures we target, the latency gap between a cache miss ($L_{miss}$) and a hit ($L_{hit}$) is typically several orders of magnitude (Li et al., 2024; Charyyev et al., 2020). As the average query latency can be modeled as $L_{avg,i}\approx HR_{i}\cdot L_{hit}+(1-HR_{i})\cdot L_{miss}$, maximizing the throughput-weighted sum of hit rates is the most direct way to minimize system-wide latency.
This leads to the following optimization problem: At each time step $t$, our goal is to find an allocation plan $a(t)=\{a_{1}(t),...,a_{N}(t)\}$ that solves the following constrained optimization problem:
(1) $\max_{a(t)}$ $\sum_{i=1}^{K}\left(\textit{ops}_{i}(t)\times HR_{i}\left(a_{i}(t)\right)\right)$
subject to $\sum_{i=1}^{K}a_{i}(t)=C_{\text{total}},$
$a_{i}(t)\geq\ell_{i},\quad\forall i\in\{1,.,K\}.$
where $C_{\text{total}}$ is the total available cache and $\ell_{i}$ is a configurable, per-tenant minimum allocation requirement to satisfy potential SLA guarantees.
Problem Hardness and Approach. The problem formulated in equation LABEL:def:problem is computationally hard. Even in its offline form, it is a variant of the Multiple-Choice Knapsack Problem (MCKP), which is NP-hard (Kleinberg and Tardos, 2006). The complexity is exacerbated in our setting, which is both online (future workloads are unknown) and black-box (performance functions are unknown a priori). Therefore, in alignment with standard practice for complex systems research, we design SAM, an autonomic control framework that employs a control-theoretic learning-based policy to find near-optimal allocations in this dynamic environment (Mao et al., 2016).
2.3. Conflicting Design Goals
The design of an effective allocation strategy is complicated by several fundamental trade-offs, which directly correspond to the challenges outlined in Section 1.
Exploitation vs. Exploration. This trade-off is central to addressing both Challenge 1 (Workload Dynamism) and Challenge 2 (Cache Pollution). An algorithm that solely exploits the current state may achieve high current throughput but risks missing emerging hotspots. Conversely, aggressive exploration by chasing naive metrics like miss count can be deceived by cache-unfriendly workloads. As detailed in Section 4, AURA resolves this by synthesizing an exploitation signal (the $\mathcal{H}$-factor) with an exploration signal (the $\mathcal{V}$-factor) via a meta-adaptive policy.
Global Throughput vs. Individual SLAs. This conflict lies at the heart of Challenge 3 (SLA Guarantees). A myopic strategy focused solely on maximizing aggregate throughput would starve low-traffic but high-priority instances. SAM’s two-pool architecture (Section 3) directly resolves this conflict by separating resources for baseline SLA guarantees from those for dynamic optimization.
Responsiveness vs. Decision Stability. This is the core trade-off for Challenge 4 (Decision Stability), especially in noisy edge environments. Overly reactive strategies are vulnerable to transient noise, causing performance oscillations that degrade sustained throughput. Conversely, excessive damping may miss genuine workload shifts. AURA’s stability-aware mechanisms, including its Saturation Confidence model and momentum-based allocator (Section 4), maintain decision consistency while preserving necessary adaptiveness.
2.4. System Model and Operating Principles
Our framework is designed based on the following principles, which we argue enhance its generality and practical relevance.
Orthogonality to Replacement Policy. Our framework focuses on dynamically sizing the cache allocated to each instance, decoupling our resource allocation logic from the internal page replacement policy (e.g., LRU, LFU) of the database engine. This is a standard separation of concerns that mirrors real-world database administration, where cache sizing is a configuration concern independent of the engine’s internal algorithms (Chou and DeWitt, 1986). This principle makes our framework general and readily deployable on top of any embedded database.
Black-Box Metric Observability. We assume that macroscopic performance metrics (e.g., ops, cache hits, misses) for each database instance are observable via an external interface. This is a practical assumption grounded in the design of modern database systems, which universally provide such observability for monitoring and administration through interfaces like ‘PRAGMA’ commands in SQLite (Hipp, 2024b) or status views in systems like MySQL and PostgreSQL (Corporation, 2024; Group, 2024). This enables our non-intrusive, external orchestrator design.
Network-Dominant Latency at the Edge. Our work specifically targets the performance challenges in modern Cloud-Edge-Device architectures (Developers, 2024b; Documentation, 2024). In such distributed systems, the primary source of user-perceived tail latency is not local disk I/O—which often occurs on fast NVMe SSDs at the edge—but the high and variable latency of network access required to fetch data from a regional data center during a cache miss. With network round-trips being typically two to three orders of magnitude slower than local storage access (Charyyev et al., 2020; Li et al., 2024), maximizing cache hits is the core optimization strategy. Other latency sources, such as lock contention, are considered orthogonal to the cache sizing problem addressed here.
3. System Overview
To address the constrained optimization problem defined in Section 2, we propose SAM, an autonomic cache management framework. The core of our stability-aware design philosophy is that principled, internal stability is the foundation for achieving both high performance and external robustness in volatile, multi-tenant environments. It moves beyond simple reactive heuristics by implementing our novel AURA algorithm, which achieves sustained high performance through a dual-factor model that balances historical efficiency with future potential.
This section details the concrete architecture of this framework. We begin by outlining its overall structure and the core two-pool resource model, then delve into the design principles behind its autonomic control policy.
3.1. Overall Architecture
The SAM system is designed to operate as an external, non-intrusive coordinator for a set of co-located, black-box database instances. As illustrated in Figure 1, the architecture is composed of two primary conceptual layers. The Resource Layer partitions the Total Host Cache ($C_{total}$) into a Fixed Pool for static baseline guarantees, and an Elastic Pool for dynamic optimization.
The System Layer contains the core control logic. A centralized SAM Autonomic Coordinator, powered by our AURA algorithm, manages the database instances residing in the Managed Environment. This architecture implements a classic feedback control loop operating in a Sense-Decide-Act cycle:
Sense: The Coordinator gathers fine-grained Performance Metrics directly from each database instance, while also perceiving global System Info from the resource layer to maintain a comprehensive view of the system state. Decide: In the decision phase, the Coordinator executes the AURA algorithm. It synthesizes the perceived state to calculate an optimal allocation plan that balances historical performance with future potential, as will be detailed in Section 4. Act: The Coordinator dispatches direct Cache Resizing Commands to each target instance, which then adjusts its memory footprint by requesting resources from or releasing them to the Elastic Pool.
This interaction is mediated by a set of simple, abstract interfaces. The Coordinator assumes only that a target instance provides a way to get metrics (e.g., ‘get_metrics()’) and a way to configure its cache size (e.g., ‘set_cache_size()’). This separation of concerns enhances the framework’s generality, allowing it to be adapted to any embedded database that exposes similar observability and configurability.
Figure 1. The architecture of the SAM system, illustrating its two-pool resource model and Sense-Act control loop. A central SAM Autonomic Coordinator perceives metrics from the database instances (Sense) and issues direct resizing commands back to them (Act), managing a partitioned Fixed Pool and Elastic Pool.
3.2. The Two-Pool Cache Architecture
A key architectural principle of our framework is the two-pool cache architecture, designed to resolve the inherent conflict between guaranteeing individual SLAs and optimizing for global performance. It operates by partitioning the total cache budget, $C_{total}$, into two distinct, non-overlapping pools. This decomposition is justified by the principle of separation of concerns: the problem of satisfying hard performance constraints is architecturally separated from the problem of opportunistic, best-effort optimization.
Fixed Pool ($C_{fixed}$). This pool embodies our architectural principle that deterministic performance baselines for critical instances should be guaranteed through static allocation, not pursued as a goal within the dynamic optimization loop. This guarantee is realized through a two-step, policy-driven allocation process. First, an administrator defines the total size of the fixed pool, $C_{fixed}$, based on overall system service-level objectives. Second, this pool is apportioned among instances ($d_{j}\in D_{p}$) proportionally to their static, pre-defined ,manually configured base_priority values. This static allocation, $a_{j}^{\text{fixed}}$, insulates critical instances from the resource volatility of the wider system. Critically, this separation of concerns allows the manager for the remaining Elastic Pool to focus single-mindedly on maximizing overall system utility, without being constrained by the safety requirements of individual tenants.
Elastic Pool ($C_{elastic}$). This pool contains all remaining cache resources ($C_{elastic}=C_{total}-C_{fixed}$) and is dedicated to maximizing the aggregate system utility, as defined by our objective function. The resources in this pool are entirely fluid and are dynamically reallocated by our core AURA algorithm, which will be introduced in the next section, in each tuning interval among all participating database instances. Finally, total allocation for any instance $d_{i}$ is the sum of its parts: $a_{i}^{\text{total}}(t)=a_{i}^{\text{fixed}}+a_{i}^{\text{elastic}}(t)$.
3.3. The SAM Coordinator and the AURA Control Policy
The autonomic control policy of the SAM framework resides in the control policy used by the Coordinator to manage the Elastic Pool. A simple policy, such as a greedy heuristic, is insufficient to navigate the conflicting challenges of dynamism, pollution, and stability defined in Section 2. Therefore, a principled, multi-factor control policy is required.
The design of our control policy is guided by two foundational principles. First, it is based on a dual-factor utility paradigm that resolves the exploitation-exploration trade-off by evaluating each instance along two orthogonal dimensions: its proven historical contribution (exploitation) and its marginal potential for future improvement (exploration). Second, it treats stability as a first-class citizen. Here, we distinguish between decision stability—maintaining consistent optimization directions despite noisy observations—and mere response damping that would slow adaptation. AURA employs control-theoretic mechanisms to achieve the former, dampening internal oscillations while ensuring robust adaptations against external disturbances.
Our novel AURA algorithm, whose detailed specification is the subject of the next section, is the concrete implementation of this control policy. By synergistically synthesizing these principles, AURA provides the autonomic control policy that allows SAM to make stable, robust, and high-performance decisions, overcoming the limitations of simpler baselines.
Figure 2. The two-level architecture of the AURA algorithm, showing: (a) the macro-level state machine that transitions between a fast-path Localized Optimization and an infrequent Global Scan to ensure scalability, and (b) the micro-level decision pipeline executed within the fast path.
4. The AURA Algorithm
The design principles articulated in the previous section culminate in the AURA algorithm, the principled, dual-factor control core of our SAM framework. AURA is architected as a two-level, adaptive state management algorithm that implements our stability-aware control policy while maintaining computational efficiency suitable for large-scale deployments. Figure 2 provides a complete visual schematic of this architecture.
4.1. Architectural Principles of the AURA Algorithm
As established in Section 2, the online, black-box nature of our problem makes finding a provably optimal solution intractable. Therefore, our approach is to design an autonomic control algorithm, AURA. AURA navigates the fundamental trade-off between exploitation and exploration to achieve the stability and robustness necessary for sustained high performance. It materializes this principle by synthesizing two orthogonal signals—a database’s proven, historical efficiency and its potential for future performance gains—allowing it to make sound decisions in complex, dynamic environments.
At a macro-level, as shown in Figure 2, AURA employs an Adaptive Active Set (AAS) framework, which operates as a two-state machine. The algorithm spends the majority of its cycles in the computationally efficient State 1 (Localized Optimization), focusing only on a small subset of tenants ($k\ll K$) to make rapid, high-quality decisions. It transitions to the more comprehensive State 2 (Global Scan) only when a strategic opportunity for improvement is detected (i.e., the active set converges). This design decouples decision latency from the system scale, providing excellent scalability while maintaining decision quality.
At a micro-level, Figure 2 details the multi-stage decision pipeline executed within the fast-path (State 1). This pipeline is centered on the dual-factor scoring model through Signal Synthesis. It evaluates each instance using a dual-factor scoring system that balances the historical efficiency ($\mathcal{H}$-factor) with the future potential ($\mathcal{V}$-factor). The recommended plan generated by this process is then passed through a final Actuation Filtering stage. If deemed significant, the plan is translated into action by a stability-enforcing actuator—our stateful, momentum-based gradient allocator—which acts as a “shock absorber” to ensure all reallocations are smooth and deliberate.
In the following subsections, we will dissect each of these components in detail, from the core factor calculations to the state transition logic.
4.2. Algorithm Specification
We now provide the complete specification of the AURA algorithm, translating the high-level architecture from the previous section into a concrete implementation. We begin by laying out the necessary data structures and state variables, followed by a detailed walkthrough of the main decision workflow and its innovative components as visually summarized in Figure 2.
4.2.1. Data Structures and State.
The AURA algorithm maintains state at two levels. The Global Coordinator State ($S_{g}$) includes: (i) Active Set State, which contains the current_active_set of tenants being optimized and cached candidate pools (top_candidates, bottom_candidates); (ii) Convergence State, such as a score_impro-
vement_window
and an inactivity_timer for stagnation detection; and (iii) Global Policy State, primarily the smoothed adaptive weight, alpha_prev. At the per-instance level, the Per-Instance State ($d_{i}$) tracks: (a) Metric State, including smoothed inputs like ema_ops_slow; (b) $\mathcal{V}$-Factor State, such as last_allocation_pages to compute gradients; and (c) Stability Control State, which features the crucial saturation_confidence score.
4.2.2. The Main Decision Workflow.
AURA’s core logic is architected as an adaptive, two-state workflow that dynamically focuses computation where it is most needed, ensuring high-quality decisions with excellent computational efficiency. The high-level process is visually summarized in Figure 2 and specified in Algorithm 1.
The algorithm’s autonomic control policy employs an adaptive two-state architecture for computational efficiency. At the start of each cycle, a key decision is made (Algorithm 1, Line 1): should it enter the infrequent, comprehensive Global Scan state, or continue in the lightweight, default Localized Optimization state? This decision is governed by a convergence metric that determines when the optimization potential of the current active set has been exhausted.
If a global scan is triggered (e.g., on a cold start or after convergence), the system performs a highly efficient, linear-time pass using a two-way heap filter to gather fresh candidates (Line 1). After passing an equilibrium check, a knee-point detection algorithm robustly determines the optimal size for a new active set (Line 1).
However, the majority of cycles are spent in the fast-path Localized Optimization state. Here, the computationally intensive scoring and allocation adjustments are confined to the small, $O(k)$ active set via the OptimizeInActiveSet function (Line 1). This two-state architecture is the key to AURA’s ability to provide high-quality decisions at a near-constant low cost, regardless of system scale. The core components used within the OptimizeInActiveSet function—the dual-factor scoring model and its supporting mechanisms—are detailed below.
Input: Set of all instances $D$ (size $K$), Global State $S_{g}$, Current Allocations $C_{t-1}$
Output: Final, executable allocation plan $C_{t}$
1 Function AURA_RunDecisionCycle:
[…]
 // Stage 1: Refresh active set if necessary
[…]
    // Slow Path logic to get a new active set
4    $TopK,BottomK\leftarrow\operatorname{TwoWayHeapFilter}(D,S_{g},k_{\max})$
5    if $\operatorname{IsSystemInEquilibrium}(TopK,BottomK)$ then
6       return $C_{t-1}$
7
8      end if
9    $k_{\text{demand}}\leftarrow\operatorname{FindKneePoint}(TopK)$
10    $S_{g}.\text{ActiveSet}\leftarrow\operatorname{ComposeSet}(TopK,BottomK,k_{\text{demand}})$
11
12   end if
13
 // Stage 2: Always perform optimization on the current (possibly new) active set
14 $C_{t}\leftarrow\textnormal{{OptimizeInActiveSet(}}\textnormal{\emph{$S_{g,t},C_{t-1}$}}\textnormal{{)}}$
15 if $C_{t}\neq C_{t-1}$ then
[…]
17
18   end if
19 return $C_{t}$
20
21
22
Algorithm 1 The AURA Dual-Factor Control Algorithm
4.2.3. Core Decision Components.
The Horizontal Factor ($\mathcal{H}$). The Horizontal Factor ($\mathcal{H}$) provides a stable, long-term measure of a tenant’s effective contribution, acting as the robust “exploitation” anchor in our system. Its design ensures both metric stability and fairness across tenants through three key architectural choices: First, its inputs are derived from a slow-speed EMA ($\lambda_{slow}=0.1$) to filter out transient performance noise and reflect only sustained contribution. Second, it employs a linear min-max normalization on throughput, placing all tenants onto a common ‘[0,1]’ scale to prevent a single high-TPS instance from dominating the score. Finally, its multiplicative form, $\mathcal{H}_{i}=\texttt{norm\_ops}_{i}\times\texttt{hr\_norm}_{i}$, acts as a crucial efficiency filter, ensuring that high throughput is only rewarded when accompanied by a high hit rate, thus inherently penalizing cache-polluting workloads.
The Vertical Factor ($\mathcal{V}$). The $\mathcal{V}$-factor provides a forward-looking, “exploratory” estimate of marginal gain. As the raw gradient, $\mathcal{V}_{raw}=\Delta HR/\Delta\texttt{allocation}$, is extremely susceptible to noise, AURA computes the final $\mathcal{V}$-factor via a principled, multi-stage signal processing pipeline. First, it applies an asymmetric EMA smoothing mechanism. This “fast-up, slow-down” policy uses a higher learning rate for positive trends (gains) and a lower rate for negative ones, allowing the system to react quickly to opportunities while cautiously evaluating performance dips. Second, the smoothed signal is passed through a robust normalization step, which scales all values based on the system-wide 90th percentile (p90), producing a statistically meaningful and outlier-resistant measure. Finally, the processed signal undergoes an additional stability layer through the Saturation Confidence model, which replaces a brittle, binary freeze with progressive influence adjustment, providing a second line of defense against erratic behavior while maintaining adaptiveness. While the asymmetric EMA handles general measurement noise, the Saturation Confidence model specifically addresses the unique instability patterns near cache saturation points.
The Meta-Adaptive Policy ($\alpha_{t}$). To dynamically balance the $\mathcal{H}$-factor (exploitation) and $\mathcal{V}$-factor (exploration), AURA employs a meta-adaptive policy that computes the weight $\alpha_{t}$. This policy acts as a system-wide “confidence dial”, driven by the insight that the statistical variance of $\mathcal{V}$-factors across the active set serves as a robust proxy for environmental uncertainty. A normalized variance metric, $\kappa$, is calculated; a high $\kappa$ signifies a volatile state with contradictory signals, pushing the system to be more conservative by increasing the weight on the stable $\mathcal{H}$-factor. Conversely, a low $\kappa$ indicates a stable state with reliable signals, allowing the system to be more opportunistic by trusting the $\mathcal{V}$-factor. The final value is temporally smoothed to ensure the policy itself remains stable.
Stateful Actuation via Gradient Allocator. The final scores produced by our synthesis model represent an ideal distribution of resources. The score for each tenant $i$ is computed as $\text{Score}_{i}=\alpha_{t}\cdot\mathcal{H}_{i}+(1-\alpha_{t})\cdot\mathcal{V}_{i}$, where $\alpha_{t}$ is the meta-adaptive weight that increases with system uncertainty. However, to prevent disruptive, large-scale reallocations in a single cycle, AURA does not apply this target directly. Instead, the OptimizeInActiveSet function uses the scores to define a target allocation, and then feeds this target into a stateful, momentum-based gradient allocator. This allocator acts as a crucial “shock absorber” or low-pass filter. It calculates an incremental, suitably-sized step from the current allocation ($C_{t-1}$) towards the ideal target. By incorporating momentum and step-size constraints, it ensures that all resource adjustments are smooth, deliberate, and stable, translating the strategic intent into a safe, executable plan, $C_{t}$.
4.2.4. Behavioral Analysis of the Scoring Function.
To understand how the composite scoring function translates into behavior, we analyze its response to four key database archetypes. (i) The Saturated, High-Contribution Instance: This instance maintains a high score due to its large $\mathcal{H}$-factor (from high ops and hit rate). However, its $\mathcal{V}$-factor converges to zero as additional cache yields no further improvement ($\Delta HR\approx 0$), correctly preventing it from wastefully accumulating more resources. (ii) The High-Potential, Low-Resource Instance: An emerging hotspot is characterized by a high marginal gain ($\mathcal{V}$-factor), as small cache additions yield large hit rate improvements. Our normalization mechanism identifies this strong signal, boosting its score and allowing it to rapidly acquire the resources it needs. (iii) The Cache-Inefficient Workload: A “cache polluter” with high ops but persistently low hit rate receives a near-zero score. This is because its $\mathcal{H}$-factor is suppressed by the low hr_norm, and its $\mathcal{V}$-factor is zero, leading to the swift reclamation of its cache. (iv) The Quiescent Instance: An idle database has near-zero ops and no activity, causing both its $\mathcal{H}$-factor and $\mathcal{V}$-factor to be zero. Consequently, it receives no elastic cache, preserving resources for active instances.
5. Analysis
In this section, we analyze effectiveness and efficiency the proposed algorithm. For the convinience of the analysis, we first give the preliminaries and assumptions.
5.1. Preliminaries and Assumptions
We formulate elastic cache allocation as an online concave maximization over
(2) $K\;=\;\Bigl{\{}x\in\mathbb{R}_{\geq 0}^{n}\ \Big{|}\ \sum_{i=1}^{n}x_{i}=C,\;x_{i}\geq\ell_{i}\Bigr{\}},$
where $C$ is the total cache size, and $\ell_{i}$ is the lower bound for tenant $i$.
Assumption 1 (Utility / Concavity).
[…]
(3) $s_{i,t}\;=\;a\,\mathcal{H}^{th}_{i,t}+b\,\mathcal{V}_{i,t},\qquad a,b>0.$
In the implementation we use $\mathcal{H}^{impl}_{i}=\mathrm{ops}_{i}\times HR_{i}$; multiplying a concave function by a non-negative scalar preserves concavity, so the analysis still applies.
Assumption 2 (Stochastic Gradient Oracle).
At round $t$ we observe an unbiased gradient estimate $\widehat{g}_{t}=\nabla f_{t}(x_{t})+\xi_{t}$ with $\mathbb{E}[\xi_{t}]=0$ and $\|\widehat{g}_{t}\|_{2}\leq G$, where $f_{t}(x)=-U_{t}(x)$.
Concavity, Robustness, and the Role of $\mathcal{H}^{impl}$.
The $O(\sqrt{T})$ regret analysis of SAM-Core (AURA-Core) leverages Assumption 1 to view the problem through online convex optimization (OCO). Real workloads can be S-shaped or piecewise non-concave. SAM-Full does not strictly rely on global concavity: it follows locally measured marginal gains ($\mathcal{V}$) with a decaying step size and stability guards, and remains robust when concavity is violated. We relax Assumption 1 to a piecewise/approximately concave condition (App. A.1.3); the bound then degrades only by a constant factor. When non-concave segments damp $\mathcal{V}$ to avoid jitter, the monotone $\mathcal{H}^{impl}$ term continues pushing allocations across shallow valleys, preventing premature stagnation. Empirically (Sec. 6.6), SAM stays within 5% of an oracle allocation even when subjected to demanding, high-contention workloads generated by the industry-standard YCSB framework. To create a challenging and non-ideal environment that violates the theoretical concavity assumption, we specifically configured YCSB for high concurrency and large record sizes, thereby inducing significant memory pressure.
Bridging Theory and Implementation.
The model above is an idealized abstraction used to prove properties of a simplified baseline, SAM-Core (AURA-Core), which directly balances $\mathcal{H}^{th}$ and $\mathcal{V}$. The full SAM algorithm adds practical heuristics for stability and scalability. Section 6 shows empirically that these additions improve performance over both SAM-Core and external baselines.
5.2. Optimality of SAM-Core (AURA-Core)
SAM-Core is our heuristic-free instantiation and reduces to Online Frank-Wolfe (OFW) (Allen-Zhu and Hazan, 2016). Given a stochastic gradient $\widehat{g}_{t}$ of the convex loss $f_{t}$ at round $t$:
(4) $y_{t}=\arg\min_{y\in K}\langle\widehat{g}_{t},y\rangle,\qquad x_{t+1}=(1-\eta_{t})x_{t}+\eta_{t}y_{t}.$
Theorem 5.1 (Static Regret).
Under Assumptions 1–2 and $\eta_{t}=2/(t+2)$,
$\mathbb{E}[\mathrm{Reg}_{T}]\leq GD\sqrt{2T}+\delta T=\mathcal{O}(GD\sqrt{T}),$
with $\delta\leq L$ from page rounding. If $f_{t}$ is $\alpha$-strongly convex, the bound tightens to $\mathcal{O}\!\big{(}(G^{2}/\alpha)\log T\big{)}$. Under approximate concavity (App. A.2), constants scale by $1/\alpha$ but rates are unchanged.
Proof sketch. The OFW rate follows Hazan (Hazan, 2016). Integer rounding adds at most $L$ loss per round. Stability guards perturb by $O(\eta_{t})$ and thus do not change the leading term.
5.3. Stability of SAM-Full
We study the full SAM algorithm with momentum and step-size decay. Let $\Delta_{t}=\|x_{t}-x_{t-1}\|_{1}$ denote the per-round allocation change. Our goal is to bound $\Delta_{t}$ and its cumulative sum.
Design Invariants.
The added stability modules enforce: (I1) bounded, decaying steps $\|\Delta_{t}\|_{1}\leq G\,\eta_{t}$ with $\eta_{t}=\Theta(t^{-1})$; (I2) suppression of updates when the V-signal is statistically untrustworthy (Saturation Confidence); (I3) directional inertia via momentum to avoid noisy oscillations.
These yield the recursion
(5) $\Delta_{t}\;\leq\;(1-\beta)\,\Delta_{t-1}\;+\;\beta G\,\eta_{t},$
with constant momentum weight $\beta\in(0,1)$.
Lemma 5.2 (Decay and Log-Bounded Variation).
With (5) and $\eta_{t}=\Theta(t^{-1})$, we have $\Delta_{t}=\mathcal{O}(t^{-1})$ and $\sum_{t=1}^{T}\Delta_{t}=\mathcal{O}(\log T)$.
Proof sketch. Unfold (5) and sum a geometric–harmonic series: $\Delta_{t}\leq(1-\beta)^{t}\Delta_{0}+\beta G\eta_{0}\!\sum_{k=1}^{t}(1-\beta)^{t-k}k^{-1}$. Summing over $t$ and using $\sum_{j\geq 0}(1-\beta)^{j}=1/\beta$ gives the $\log T$ bound; the pointwise $t^{-1}$ decay follows from dominance of the last terms. Full derivation is in Appendix A.3.
Theorem 5.3 (Regret of SAM-Full).
Since $U_{t}$ is $L$-Lipschitz, the extra loss due to stability is bounded by $L\Delta_{t}$, hence
$\sum_{t=1}^{T}(\tilde{f}_{t}-f_{t})\;\leq\;L\sum_{t=1}^{T}\Delta_{t}=\mathcal{O}(\log T).$
Combining with Theorem 5.1, the total regret is $Reg_{T}^{\textsc{SAM}}=\mathcal{O}(GD\sqrt{T})+\mathcal{O}(\log T)$; the $O(\sqrt{T})$ term dominates asymptotically.
5.4. Computational Footprint
A key goal of SAM is low decision latency at scale. We analyze the per-cycle time and memory cost of the full adaptive algorithm.
Theorem 5.4 (Per-cycle Complexity).
The full SAM algorithm has worst-case per-cycle time $O(K\log k_{\max})$, amortized time substantially below $O(K)$ (approaching $O(k)$ in steady state), and memory $O(K)$.
Proof sketch. SAM alternates between (i) a Localized Optimization phase over the active set of size $k\!\ll\!K$ (cost $O(k)$), and (ii) an infrequent Global Scan whose dominant step is a two-way heap filter over all $K$ tenants ($O(K\log k_{\max})$). Since scans occur only after many cheap local steps, the average per-cycle cost is dominated by $O(k)$. Each tenant keeps constant-size state, so memory is linear in $K$. Full derivation is in Appendix A.4.
6. Experimental Evaluation
This section presents a comprehensive experimental evaluation of our proposed system, SAM, which is powered by our novel AURA allocation algorithm. We first detail the experimental setup and baselines, then present a series of targeted experiments designed to validate SAM’s performance, stability, and scalability.
6.1. General Experimental Setup
System Environment. All experiments run on Windows 11 (Intel Core i5-12600KF, 32GB RAM) with Python 3.11.7 and APSW/SQLite 3.40.0. We stress I/O by using large records (2KB).
Figure 3. End-to-end performance on the standard hotspot-shift workload. (a) overall throughput, (b) hit rate, (c) phase-by-phase throughput. S0 and B7 are top-performing and most adaptive.
Evaluated Strategies. We compare SAM with 14 baselines covering (i) static allocations (e.g., B1, B2); (ii) ablations of SAM components (e.g., B3, B8, B10); (iii) advanced dynamic heuristics (e.g., B7, MT-LRU/B12); and (iv) a smoothed offline upper bound (B14). Detailed definitions and the implementation of B14 appear in Appendix B.1.
We omit Reinforcement Learning (RL) baselines as our target non-stationary workloads and resource-constrained edge nodes present significant challenges for the sample efficiency and continuous retraining requirements of modern RL agents. Furthermore, our Online Frank-Wolfe based SAM-Core already provides a strong theoretical regret bound, making complex RL-based approaches orthogonal to this work’s core contributions.(Mao et al., 2016, 2019)
Metrics. We evaluate both instantaneous performance and sustained performance stability through the following metrics: Scalability (CPU decision time), Productivity (total TPS, P99 latency), Efficiency (hit rate, TPS/Page), and Decision Stability (throughput stability $\sigma_{\text{TPS}}$), and Fairness (QoS for non-hotspot tenants).
Figure 4. Cache allocation dynamics during the hotspot-shift workload. S0 and B7 are stable and effective; B12 is highly unstable.
Figure 5. S0’s robustness against cache pollution. (a) The attack slightly raises the ‘ops*miss_rate’ score; (c) B7 catastrophically reallocates resources, causing (b) severe VIP degradation. S0 maintains protection.
Workload Generation. To evaluate policy adaptability, our primary experiments employ a scripted three-phase hotspot-shift workload. This workload simulates dynamic operational conditions by first establishing a baseline load, then creating a performance hotspot on a high-priority instance (Phase 2), and finally shifting this hotspot to a different, medium-priority instance (Phase 3). This design rigorously tests each policy’s ability to react to, and recover from, sudden changes in load distribution. A detailed, reproducible specification of this workload is available in Appendix B.2.
6.2. Performance on Standard Dynamic Workloads
We begin with a three-phase hotspot-shift workload to test each policy’s ability to adapt under normal dynamics before moving to adversarial cases.
End-to-End Performance. The end-to-end results on the hotspot-shift workload, summarized in Figure 3, establish SAM (S0) and the greedy heuristic (B7) as the clear top tier. Both achieve superior median throughput (34.4 and 35.0 TPS, respectively) and hit rates, driven by their adaptability to the shifting hotspot. This contrasts sharply with static policies like B2, which show minimal performance change. The allocation trajectories in Figure 4 reveal the behavior behind these results: while both S0 and B7 demonstrate stable and effective resource allocation logic by correctly reallocating cache to the new hotspot in Phase 3, other advanced policies like B12 exhibit extreme decision oscillations, resulting in degraded sustained performance despite occasional peaks.
[…]
Generalization on the TPC-C Benchmark. To validate SAM’s ability to generalize to complex OLTP workloads, we evaluate it on the industry-standard TPC-C benchmark, with results summarized in Table 2. The data reveals critical design trade-offs: while an idealized global LRU pool (B5) yields the highest throughput, its catastrophically high P99 latency makes it unsuitable for OLTP environments. The true contest is between SAM and the best-in-class greedy heuristic (B7). On this non-adversarial benchmark, SAM matches the strong baseline performance of B7 (e.g., identical median TPS) while achieving a higher mean throughput, indicating a superior ability to exploit optimization opportunities. Although B7 shows slightly better latency in this benign “peace-time” scenario, SAM’s key advantage is its proven robustness. As demonstrated in our adversarial analysis (§6.3), SAM is resilient to cache pollution attacks where B7 fails catastrophically, establishing SAM as a more reliable, production-grade solution for real-world environments.
Table 1. Comprehensive trade-off analysis of SAM against advanced dynamic strategies.
Strategy Median TPS Adaptation Lag (s) Allocation Stability ($\sigma$) Decision Overhead
SAM (ours) 34.0 38.7 17.5 O(K) 0.13ms
B11 (Regression) 33.8 53.0 41.4 O(K) ¡0.01ms
B13 (UCP) 32.2 55.6 21.8 O(P·K) 0.02ms
Table 2. Key Performance and QoS Metrics on the TPC-C Benchmark.
Strategy Median TPS Mean TPS Median P99 Latency (ms)
SAM (ours) 0.750 0.978 625.7
B2 (Static) 0.750 0.918 690.0
B7 (Greedy) 0.750 0.942 579.9
B12 (SLA-Driven) 0.500 0.606 1480.3
B5 (Global LRU) 1.050 1.073 1932.0 (unacceptable)
Table 3. Performance comparison against near-optimal baselines.
Metric SAM (ours) B5 (Global LRU) B14 (Optimal)
Phase 2 (High-Priority Hotspot)
Median TPS 33.2 19.0 33.5
Hit Rate (%) 90.9 73.5 90.7
P99 Latency (ms) 77.5 116.5 87.3
Phase 3 (Hotspot Shift)
Median TPS 35.2 21.8 35.0
Hit Rate (%) 92.2 73.6 92.5
Comparison with Near-Optimal Baselines. To precisely anchor SAM’s performance, we compare it against an idealized global LRU pool (B5) and the theoretical Hindsight Optimal policy (B14). As summarized in Table 3, the results conclusively show that SAM operates at the theoretical performance limit. Its throughput and hit rate are nearly identical to the optimal B14 baseline in both phases, while decisively outperforming the architecturally-idealized B5 policy. Crucially, SAM’s priority-aware design translates to superior Quality of Service, achieving the lowest P99 latency (77.5 ms) for the high-priority tenant—even surpassing the purely hit-rate-focused oracle policy. This establishes SAM as a near-optimal and practically superior solution.
Summary. On this standard dynamic workload, SAM proves to be a top-tier performer. It matches the throughput of the best greedy heuristic (B7) while achieving superior decision stability, outperforms advanced learners (B11) and theoretical frameworks (B13) in adaptation speed, and operates near the theoretical optimal (B14). These results demonstrate that SAM successfully achieves its design goal: sustained high performance through superior stability.
6.3. Robustness under Adversarial Workloads
Having established the strong performance of both SAM (S0) and the greedy heuristic (B7) on a standard dynamic workload, we now subject them to a more challenging, adversarial scenario. To validate resilience against (C2) cache pollution, we designed an intermittent scan attack where a low-priority tenant mimics a high-demand workload through cache-unfriendly, scan-intensive operations. This stress-test is designed to answer a critical question: can the policy distinguish between a legitimate hotspot and a deceptive “polluter”? The full failure cascade for the reactive policy is visualized in Figure 5.
Figure 6. Performance comparison of S0 and its ablation variants. The results demonstrate that removing the $\mathcal{H}$-factor (B10) or the $\mathcal{V}$-factor (B8), or making the $\mathcal{H}$-factor overly reactive (B9), all degrade overall performance. The apparent superiority of B3 is a QoS trade-off analyzed in Figure 7.
Figure 7. Quality of Service (QoS) comparison in the complex Phase 3 workload. The figure reveals the critical trade-off where B3’s “winner-take-all” strategy significantly harms non-hotspot tenants, unlike S0’s QoS protection mechanism.
Figure 8. Cache allocation dynamics of the ablation variants. The plot demonstrates that the $\mathcal{V}$-factor is critical for agility (cf. B8’s sluggishness), the $\mathcal{H}$-factor is essential for action (cf. B10’s inertness), and that the $\mathcal{H}$-factor must be conservative (cf. B9’s suboptimal and unstable behavior).
Flawed Logic and Severe Misallocation. An analysis of the resource allocation logic reveals the fundamental weakness of the shallow heuristic. The intermittent scan attack generated a mere 11.7% increase in the ‘ops*miss_rate’ “need score” for the attacker. In response, B7 irrationally increased its cache allocation to this low-value tenant by a staggering 64.5%—a quantified 5.5$\times$ overreaction. To evaluate the consequences of this flawed logic, we created a scenario where a high-priority “VIP” tenant and the attacker compete for a shared 2MB cache. SAM correctly prioritized the VIP, allocating resources at a healthy 1.4-to-1 ratio in its favor (241 vs. 172 pages). In stark contrast, B7’s logic led to a severe resource inversion, diverting the vast majority of resources to the attacker at a disastrous 1-to-4.6 ratio (a mere 73 pages for the VIP vs. 336 for the attacker).
Performance Impact and Final Verdict. This severe misallocation translated directly into degraded service quality. As shown in the time-series plot, SAM maintained the VIP tenant’s hit rate between 97-98% throughout the attack. In contrast, B7’s VIP hit rate fell from 97% to 95.5% during the sustained attack, with a sharp dip to 92%. This experiment demonstrates that B7’s simple heuristic, while effective on benign workloads, becomes a critical vulnerability under adversarial conditions. SAM’s dual-factor design—evaluating both historical contribution and marginal gain—provides the robustness necessary for sustained high performance in real-world systems where such attacks are inevitable.
6.4. Ablation Studies
To justify the design choices of our SAM algorithm that considers multiple orthogonal factors, we conducted a series of ablation studies. We evaluate the full SAM policy (henceforth S0 in figures) against four variants, each with a critical component modified or removed: B10 (no $\mathcal{H}$-factor), B8 (no $\mathcal{V}$-factor), B3 (no Fixed Pool), and B9 (a fast, reactive $\mathcal{H}$-factor). The results are shown in Figures 6, 7, and 8.
Essentiality and Nature of the $\mathcal{H}$ and $\mathcal{V}$ Factors. The results unequivocally demonstrate that both the $\mathcal{H}$ and $\mathcal{V}$ factors are essential, and that the $\mathcal{H}$-factor must be deliberately conservative. As shown in the overall performance summary (Figure 6a), removing either factor is detrimental. B10 (no $\mathcal{H}$-factor) is the worst performer with a median TPS of 32.4, while B8 (no $\mathcal{V}$-factor) also performs poorly at 33.0 TPS. The allocation dynamics in Figure 8 reveal the reasons for these failures. B10 is completely inert, failing to identify any valuable tenant. Conversely, B8 is sluggish; without the forward-looking $\mathcal{V}$-factor, it is slow to adapt to the new hotspot in Phase 2.
The case of B9, which uses a fast EMA for its $\mathcal{H}$-factor, provides a deeper insight. Despite being more reactive to changes, its overall performance (32.7 TPS) is worse than the more conservative B8. The reason is revealed in its dynamics (Figure 8): its fast-rising $\mathcal{H}$-factor causes it to prematurely converge on a suboptimal allocation of only  300 pages in Phase 2. Furthermore, at the start of Phase 3, this lingering (but quickly-formed) historical bias causes it to mistakenly allocate resources to the now-cold high-priority database. This proves that the $\mathcal{H}$-factor’s role as a stable, conservative anchor (using a slow EMA) is critical, not to impede adaptation, but to ensure consistent, high-quality decisions that lead to sustained performance.
The Value of the QoS Protection Mechanism. The most nuanced result comes from comparing S0 with B3 (no Fixed Pool). While B3 achieves a slightly higher median TPS (34.2 vs. S0’s 34.1, Figure 6a), this minor gain masks a severe trade-off in service quality. Figure 7 provides the crucial insight. In the more complex Phase 3, to achieve a marginal +0.8% gain in hotspot hit rate (92.3% vs. S0’s 91.5%), B3 allows the median non-hotspot hit rate to drop to 96.9%. S0, in contrast, successfully protects these tenants, maintaining their hit rate at 97.9%. This represents a nearly 50% relative increase in cache misses for all other services in the system, demonstrating that B3’s “winner-take-all” strategy is detrimental to overall system health.
Summary. The ablation studies confirm that SAM’s design is a necessary synthesis of its components. The results prove three core design principles: (1) The $\mathcal{H}$-$\mathcal{V}$ dual-factor system is essential for achieving both decision stability and adaptiveness. (2) The $\mathcal{H}$-factor must be conservative and stable to properly anchor the policy’s exploration. (3) The QoS protection mechanism is critical for providing system-wide fairness in a multi-tenant environment.
6.5. Scalability and Architectural Analysis
Figure 9. Scalability analysis of SAM. The top plot shows SAM’s near-constant decision latency, which outperforms the linear scaling of the greedy heuristic (B7) at larger scales. The bottom plot reveals the reason: SAM’s two-state architectural signature, consisting of a dominant peak of ultra-fast localized optimizations and a sparse tail of rare, more expensive global scans.
To validate that SAM’s stability-aware design does not compromise scalability, we analyzed its average CPU decision time as the number of co-located databases ($K$) increased from 20 to 120. The results in Figure 9 show that SAM’s decision latency exhibits a remarkably flat, near-constant scaling profile. While the highly-optimized greedy heuristic (B7) demonstrates a clear linear increase ($O(K)$) and is faster at small scales, SAM becomes demonstrably more efficient at $K=120$. This empirically proves that our Adaptive Active Set architecture successfully decouples computational cost from the system scale.
A key observation from Figure 9(a) is SAM’s higher variance in decision time. This is not a sign of instability, but rather a direct consequence of its two-state adaptive architecture. Figure 9(b) shows the probability density for $K=120$: a sharp peak at $\approx$0.2ms (the $O(k)$ Localized Optimization mode used in most cycles) with a sparse tail extending to the 95th percentile of 0.63ms (the rare $O(K)$ Global Scans). This bimodal distribution confirms that expensive global scans occur as true exceptional events, not routine operations, keeping SAM’s amortized cost minimal despite its comprehensive optimization capabilities.
Figure 10. Theory-driven checks: (a) log–log regret of Core and Full; (b) steady-state utility ratio vs. oracle; (c) allocation jitter $\sigma_{\!\Delta}$ (lower is better).
Finally, to validate SAM’s practicality for resource-constrained environments, we deployed it on a Raspberry Pi 5. In a 20 databases scenario with an 8 seconds decision interval, SAM’s average CPU decision time was a mere 0.496ms. This constitutes only 0.0062% of the total decision budget, demonstrating that SAM is not only theoretically scalable but also lightweight enough for real-world edge deployments.
Scalability via Adaptive Active Set. To ensure scalability, SAM employs an Adaptive Active Set (AAS) framework, which confines computation to a small subset of tenants in most cycles. This mechanism is the key to SAM’s near-constant decision latency (as shown in §6.8), but it can introduce a transient delay in reacting to nascent hotspots. For this reason, and because its overhead outweighs its benefits for small populations, the AAS is explicitly disabled in all experiments with fewer than 10 tenants, including our core end-to-end evaluations in §6.2 and §6.3.
6.6. Empirical Checks for Theoretical Guarantees
This subsection empirically validates the key theoretical claims made in §5. We specifically aim to demonstrate: (i) that SAM-Core achieves the optimal $\mathcal{O}(\sqrt{T})$ static regret bound for an online algorithm; and (ii) that the full SAM algorithm’s stability mechanisms successfully bound the allocation change, leading to a regret that is asymptotically identical to its theoretical core.
Method. We rerun the 10-tenant workload (Phases 1–3) and log per-round loss $f_{t}$, oracle loss $f_{t}^{\star}$, and allocation change $\Delta_{t}$. Regret is plotted on a log–log scale; jitter is summarised by the standard deviation $\sigma_{\!\Delta}$ over the steady-state window (last 10% rounds). Each curve averages 5 runs; bands show 95% CI.
[…]
7. Related Work
Our work on stability-aware, adaptive cache management for multi-tenant embedded databases, which uniquely treats decision stability as a prerequisite for sustained high performance, intersects with several established research areas. We position our contributions with respect to prior work in database cache management, resource allocation in multi-tenant systems, and the application of online learning and control theory to resource scheduling.
Database Cache Management. The problem of optimizing cache performance within a single database instance is well-studied. Research in this area can be broadly categorized into page replacement policies and buffer pool sizing.
Page replacement algorithms determine which page to evict when the cache is full. While the classic Least Recently Used (LRU) policy is simple, it is notoriously vulnerable to workloads with poor temporal locality. To address this, numerous advanced policies have been proposed, such as X3 (Johnson and Shasha, 1994), LIRS (Jiang and Zhang, 2002), and self-tuning algorithms like ARC (Megiddo and Modha, 2003) and CAR (Bansal et al., 2004). Another line of work focuses on dynamically tuning the buffer pool size for a single database, as pioneered by works like DBMIN (Chou and DeWitt, 1986).
Our work is orthogonal to these internal optimization techniques. We do not propose a new page replacement policy. Instead, SAM operates at a higher level of abstraction, addressing the inter-instance cache partitioning problem: how to dynamically allocate a total cache budget among multiple, independent, “black-box” database instances.
Resource Management in Multi-Tenant and Cloud Environments. Managing shared resources to provide performance isolation is a central challenge in multi-tenant systems. Seminal work like SQLVM (Narasayya et al., 2013) and large-scale cloud databases like Amazon Aurora (Verbitski et al., 2017) have developed techniques for resource management in monolithic, white-box architectures. Other research has focused on mitigating the “noisy neighbor” problem for CPU (Das et al., 2013) and I/O (Gulati et al., 2010).
In stark contrast, SAM is specifically designed for the unique challenges of edge computing environments, where co-located services are often fully independent, heterogeneous, black-box processes. Our framework acts as an external, non-intrusive coordinator, providing intelligent resource management without requiring any modification to the database engines themselves. The design philosophy of our Two-Pool Architecture is inspired by mature paradigms from system-level resource management like Kubernetes (Burns and Oppenheimer, 2016) resource ‘requests’ and ‘limits’, and memory ballooning in hypervisors (Waldspurger, 2002).
Online Learning and Control for Resource Management. Our work builds upon Online Convex Optimization (OCO) (Hazan, 2016) for theoretical foundations and relates to utility-based cache partitioning like UCP (Qureshi and Patt, 2006). However, SAM advances this paradigm with a novel dual-factor model synthesizing historical efficiency ($\mathcal{H}$-factor) and marginal gain ($\mathcal{V}$-factor), providing a robust framework for the exploitation-exploration trade-off (Bubeck et al., 2012). Our stability mechanisms draw from control theory (Åström and Murray, 2021). SAM’s novelty lies in recognizing that stability is not a constraint but an enabler of sustained performance, and in the synergistic synthesis of online learning and control theory principles to achieve this stability-aware design.
8. Conclusion
In this paper, we introduced SAM, a stability-aware cache manager that addresses the critical challenge of resource contention in multi-tenant embedded databases. Our key insight is that sustained high performance requires principled stability, not reactive adaptation. SAM’s core, the novel AURA algorithm, implements this philosophy by balancing historical efficiency ($\mathcal{H}$-factor) and forward-looking marginal gain ($\mathcal{V}$-factor). Built on a Two-Pool model to decouple guarantees from optimization, SAM proved superior in extensive evaluation against 14 baselines, achieving near-optimal throughput while uniquely resisting adversarial workloads. These results validate our thesis: decision stability is the foundation for sustained high performance in real-world systems. Future work could extend these principles to proactive allocation, energy efficiency, or multi-resource co-management.
Appendix A Theoretical Analysis and Proofs
A.1. Additional Details for Section 5.1
A.1.1. Full Assumptions, Notation, and Relaxations
Feasible Set.
$K=\{x\in\mathbb{R}^{n}_{\geq 0}\mid\sum_{i}x_{i}=C,\ x_{i}\geq\ell_{i}\}$ is a compact convex polytope with Euclidean diameter $\operatorname{diam}_{2}(K)=D$.
Assumption 1 (Utility / Concavity, full statement).
For each tenant $i$, $HR_{i}:[0,C]\to[0,1]$ is concave, non-decreasing, and $L$-Lipschitz. Define
$\mathcal{H}^{th}_{i}(x)=HR_{i}(x),\qquad\mathcal{V}_{i}(x)=\frac{\partial HR_{i}(x)}{\partial x}.$
The score is $s_{i,t}=a\,\mathcal{H}^{th}_{i,t}+b\,\mathcal{V}_{i,t}$ with $a,b>0$ (we set $a=b=0.5$ in experiments; $b$ absorbs the scale factor $\lambda$).
Assumption 1’ (Approximate / Piecewise Concavity).
There exists $\alpha\in(0,1]$ such that for any $x,y\in[0,C]$ and $\lambda\in[0,1]$,
$HR_{i}(\lambda x+(1-\lambda)y)\ \geq\ \alpha\bigl{[}\lambda HR_{i}(x)+(1-\lambda)HR_{i}(y)\bigr{]}.$
Equivalently, $HR_{i}$ is dominated by a concave envelope within factor $1/\alpha$. Under Assumption 1’, the $O(\sqrt{T})$ regret bound scales by $1/\alpha$.
Implementation $H$-factor.
In deployment we use $\mathcal{H}^{impl}_{i}=\mathrm{ops}_{i,t}\cdot HR_{i}(x_{i,t})$. Since $\mathrm{ops}_{i,t}\geq 0$ is exogenous to $x_{i,t}$ within round $t$, $f_{t,i}(x)=\mathrm{ops}_{i,t}\cdot(1-HR_{i}(x))$ remains convex in $x$, and multiplying by a non-negative scalar preserves concavity of $HR_{i}(\cdot)$.
Assumption 2 (Stochastic Gradient Oracle, full).
We observe $\widehat{g}_{t}=\nabla f_{t}(x_{t})+\xi_{t}$, where $\mathbb{E}[\xi_{t}\mid\mathcal{F}_{t-1}]=0$, $\|\widehat{g}_{t}\|_{2}\leq G$, and $\{f_{t}\}$ may vary adversarially but each $f_{t}$ is convex in $x$ (concave in utility).
Additional Symbols.
Table 4 summarizes symbols used in §5.
Table 4. Notation for the theoretical analysis (expanded).
Symbol Description
$K$ Feasible set; $\operatorname{diam}_{2}(K)=D$
$C$ Total cache capacity
$\ell_{i}$ Lower bound for tenant $i$
$w_{t,i}$ Request volume of tenant $i$ at round $t$
$U_{t}(x)$ Utility at round $t$; $f_{t}(x)=-U_{t}(x)$ is the loss
$\widehat{g}_{t}$ Stochastic gradient, $\|\widehat{g}_{t}\|_{2}\leq G$
$Reg_{T}$ Static regret $\sum_{t=1}^{T}(f_{t}(x_{t})-\min_{x\in K}f_{t}(x))$
$\Delta_{t}$ Allocation change $\|x_{t}-x_{t-1}\|_{1}$
$\eta_{t}$ Step size, $\eta_{t}\leq\eta_{0}t^{-p}$ with $p\in(0.5,1]$
$\alpha$ Approximate concavity factor (Assumption 1’)
A.1.2. Why Concavity and How We Relax It
Concavity enables an OCO treatment and the Frank–Wolfe style update in SAM-Core, yielding $O(\sqrt{T})$ static regret. When $HR_{i}$ is only approximately concave (Assump. 1’), we get the same bound up to a factor $1/\alpha$. For highly non-concave segments, SAM-Full’s stability guards (momentum, significance gating, saturation) bound total variation $\sum_{t}\Delta_{t}=O(\log T)$, ensuring the algorithm does not oscillate.
[…]
A.1.3. Link to the Implementation
SAM-Core (AURA-Core) directly uses $(\mathcal{H}^{th},\mathcal{V})$ without heuristic guards; SAM-Full adds them for stability and scalability. Because each guard perturbs updates by $O(\eta_{t})$, the $O(\sqrt{T})$ term remains dominant; the added variation contributes only a constant.
Lemma A.1 (Constant Additive Gap).
Let $x_{t}^{\text{core}}$ be the iterate of SAM-Core and $x_{t}$ that of SAM-Full. Assume the stability guards add a perturbation $\phi_{t}$ with $\|\phi_{t}\|_{2}\leq c\,\eta_{t}$ and $\eta_{t}\leq\eta_{0}t^{-p}$, $p>0.5$. Then
$\sum_{t=1}^{T}\bigl{(}f_{t}(x_{t})-f_{t}(x_{t}^{\text{core}})\bigr{)}\leq Gc\sum_{t=1}^{\infty}\eta_{t}=O(1).$
Hence the regret of SAM-Full differs from SAM-Core’s by at most a constant.
Proof sketch. Triangle inequality gives $|f_{t}(x_{t})-f_{t}(x_{t}^{\text{core}})|\leq G\|x_{t}-x_{t}^{\text{core}}\|_{2}$. The one-step deviation is bounded by $\|x_{t}-x_{t}^{\text{core}}\|_{2}\leq\|\phi_{t}\|_{2}\leq c\eta_{t}$. Summing over $t$ and using $\sum_{t}\eta_{t}<\infty$ (since $p>0.5$) yields the claim. ∎
A.2. Details for Section 5.2: Regret of SAM-Core
OFW Update on a Simplex.
Given $\widehat{g}_{t}$, solve $y_{t}=\arg\min_{y\in K}\langle\widehat{g}_{t},y\rangle$. Since $K$ is a simplex with lower bounds $\ell_{i}$, the solution assigns all mass to the coordinate with smallest gradient component, then adjusts proportionally to respect $\sum_{i}y_{i}=C$ and $y_{i}\geq\ell_{i}$. The update $x_{t+1}=(1-\eta_{t})x_{t}+\eta_{t}y_{t}$ is a Frank–Wolfe step.
Static Regret Bound (Full Statement).
Let $x^{\star}=\arg\min_{x\in K}\sum_{t=1}^{T}f_{t}(x)$. Denote the diameter $D=\operatorname{diam}_{2}(K)$. With $\eta_{t}=2/(t+2)$ and $\|\widehat{g}_{t}\|_{2}\leq G$, we have
$\mathbb{E}\big{[}\mathrm{Reg}_{T}\big{]}=\mathbb{E}\Big{[}\sum_{t=1}^{T}f_{t}(x_{t})-f_{t}(x^{\star})\Big{]}\leq GD\sqrt{2T}+\delta T.$
The proof follows Hazan’s analysis of OFW (Hazan, 2016). The rounding term $\delta T$ arises because we store pages as integers; per-round utility loss is at most $L$ due to $L$-Lipschitz continuity, hence $\delta\leq L$.
Strong Convexity Case.
If each $f_{t}$ is $\alpha$-strongly convex, applying the standard OFW result yields
$\mathbb{E}[\mathrm{Reg}_{T}]=\mathcal{O}\!\big{(}(G^{2}/\alpha)\log T\big{)}.$
Approximate Concavity.
Under Assumption 1’ (App. A.1.1), the regret bounds scale by at most $1/\alpha$; the rates $\sqrt{T}$ and $\log T$ remain unchanged.
Rounding Error Bound.
Let $\bar{x}_{t}$ be the fractional iterate before rounding and $x_{t}$ the rounded version. Then $\|x_{t}-\bar{x}_{t}\|_{1}\leq n$ pages, but only 1 page per tenant changes; hence per-round loss increment $\leq L$ by Lipschitzness. Summing over $T$ gives $\delta T$.
Connection to SAM-Full.
Let $x^{\text{core}}_{t}$ be SAM-Core’s iterate and $x_{t}$ that of SAM-Full. Stability guards add a perturbation $\phi_{t}$ with $\|\phi_{t}\|_{2}\leq c\eta_{t}$ (App. A.1.3). Then
$\sum_{t=1}^{T}\bigl{(}f_{t}(x_{t})-f_{t}(x^{\text{core}}_{t})\bigr{)}\leq Gc\sum_{t=1}^{\infty}\eta_{t}=O(1),$
so SAM-Full’s regret differs by only a constant term.
A.3. Derivation of the Stability Bounds (Details for § 5.3)
From Guards to Recursion.
Let $x_{t}^{\text{core}}$ be the OFW update before guards, and $x_{t}$ the final allocation. The guards (momentum, saturation, significance gating) modify the step by $\phi_{t}$ with $\|\phi_{t}\|_{1}\leq G\eta_{t}$. Momentum yields
$x_{t}=(1-\beta)\,x_{t-1}+\beta\,\tilde{x}_{t},\quad\tilde{x}_{t}=x_{t-1}-\eta_{t}\widehat{g}_{t}+\phi_{t},$
hence
$\Delta_{t}=\|x_{t}-x_{t-1}\|_{1}$ $\leq(1-\beta)\|x_{t-1}-x_{t-2}\|_{1}+\beta\|\eta_{t}\widehat{g}_{t}-\phi_{t}\|_{1}$
$\leq(1-\beta)\Delta_{t-1}+\beta G\eta_{t},$
establishing (5).
Proof of Lemma 5.2.
Iterating the recursion,
$\Delta_{t}\leq(1-\beta)^{t}\Delta_{0}+\beta G\sum_{k=1}^{t}(1-\beta)^{t-k}\eta_{k}.$
With $\eta_{k}=\eta_{0}k^{-p}$ and $p=1$ (our implementation) we bound
$\sum_{k=1}^{t}(1-\beta)^{t-k}k^{-1}\leq\sum_{j=0}^{\infty}(1-\beta)^{j}(t-j)^{-1}\leq\frac{1}{\beta}\cdot\frac{1}{t}+\frac{1}{\beta}\sum_{m=1}^{t-1}\frac{1}{m}$
$=\mathcal{O}\bigl{(}\tfrac{1}{t}\bigr{)}+\mathcal{O}(\log t),$
which yields $\Delta_{t}=\mathcal{O}(t^{-1})$ and $\sum_{t=1}^{T}\Delta_{t}=\mathcal{O}(\log T)$.
theorem 5.3.
Because $U_{t}$ is $L$-Lipschitz w.r.t. $\|\cdot\|_{1}$, $\tilde{f}_{t}-f_{t}\leq L\Delta_{t}$. Summation gives an $O(\log T)$ additive term. Hence $Reg_{T}^{\textsc{SAM}}=Reg_{T}^{\textsc{Core}}+O(\log T)$, dominated by $O(\sqrt{T})$.
A.4. Derivation of Computational Complexity (Details for § 5.4)
Two-Phase Execution Model.
Let $K$ be the total number of tenants, $k$ the active set size, and $k_{\max}$ the cap used in global scans.
  • Localized Optimization: Runs every cycle. Gradient allocation, guard checks, and score updates touch only the $k$ active tenants. Each operation is $O(k)$.
  • Global Scan: Triggered when the local phase converges or a significance test fires. We identify top/bottom $k_{\max}$ candidates via a two-way heap in $O(K\log k_{\max})$. Because $k_{\max}$ is a small constant, this is effectively linear in $K$.
Amortized Time.
Suppose one global scan is followed by $m$ local cycles. The total cost is
$T_{\text{total}}=O(K\log k_{\max})+m\cdot O(k).$
Hence the amortized per-cycle cost is
$\bar{T}=\frac{T_{\text{total}}}{m+1}=O\!\Big{(}\frac{K\log k_{\max}}{m}\Big{)}+O(k),$
which approaches $O(k)$ as $m$ grows (empirically $m\gg 1$ in steady state).
Memory.
We keep a constant-size record per tenant (scores, last allocation, guard state), yielding $O(K)$ total memory. .
Trigger Frequency.
Global scans are infrequent and triggered only after many cheap local cycles, so the amortized cost approaches $O(k)$. (See Sec. 6.5 for empirical latency.)
Appendix B Additional Experimental Results
B.1. Baselines
Table 5. Summary of Evaluated Strategies
Strategy ID Core Idea and Rationale
SAM (Ours) Our proposed system, implementing the AURA algorithm. It combines a dual-factor (H/V) scoring model with an Adaptive Active Set framework for stability and scalability.
Static Baselines
B1_STATIC_AVERAGE Equal cache allocation; a simple fairness baseline.
B2_FIXED_PRIORITY Priority-based static allocation; a common real-world heuristic.
B6_DATASIZE_PROP Data-size-based static allocation.
Ablation Studies
B3_PURE_ELASTIC SAM with the fixed pool disabled; tests the value of the QoS mechanism.
B8_EFFICIENCY_ONLY SAM with the V-factor disabled; tests the value of exploration.
B10_POTENTIAL_ONLY SAM with the H-factor disabled; tests the value of historical stability.
Advanced Dynamic Baselines
B4_INDIVIDUAL_OPT Selfish, uncoordinated greedy allocation.
B5_GLOBAL_LRU Simulates an ideal, unified global LRU cache pool.
B7_DYNAMIC_NEED Allocates cache based on the popular ‘ops * miss_rate‘ heuristic.
B11_REGRESSION Allocates based on regression curve fitting of cache performance.
B12_MT_LRU An academic State-of-the-Art, SLA-driven policy.
B13_UCP A policy based on the marginal utility principles of UCP.
Theoretical Upper Bound
B14_HINDSIGHT_OPT The offline, per-phase optimal static allocation, serving as a theoretical ceiling.
We list B1–B14, their objective/heuristic, and required inputs.
B14_HINDSIGHT_OPT (Smoothed Offline Optimal)
This strategy serves as a practical theoretical upper bound for each workload phase and is not a real-time algorithm. To compute its allocation, we first employed an extensive offline profiling process. For each distinct workload phase, we empirically measured the hit-rate curve, $HR_{i}(c)$, for every database $i$ by running the trace with a wide range of discrete cache allocations $c$. With this hindsight knowledge, the problem of finding the allocation that maximizes our objective function (as defined in Section 2) is framed as a Multiple-Choice Knapsack Problem (MCKP). In this model, each database represents an item group, discrete cache chunks are the items within each group, the size of a chunk is its “weight”, and the resulting effective throughput ($\textit{ops}_{i}\times HR_{i}(c)$) is its “value”. We then solved this NP-hard problem using a dedicated integer programming solver to find the optimal static allocation for each phase.
B.2. Workload Specifications
The primary workload used for end-to-end performance evaluation (e.g., in Section 6.2) is a scripted 660-second, three-phase scenario designed to test policy adaptability. The target environment consists of multiple database instances with varying base priorities (e.g., High, Medium, Low).
  1. (1)
    Phase 1: Baseline (60s):
    • Action: All database instances are subjected to a stable baseline workload with a nominal request rate (1x).
    • Purpose: Allows all policies to warm up their caches and reach a steady state, establishing a fair performance baseline for comparison.
  2. (2)
    Phase 2: High-Priority Hotspot (300s):
    • Action: The request rate to the designated ‘db_high_prio’ instance is increased by a factor of 30x. All other instances remain at the 1x baseline load.
    • Purpose: To simulate a sudden surge in demand for a critical service and evaluate the policy’s ability to swiftly identify and allocate resources to a new hotspot.
  3. (3)
    Phase 3: Hotspot Shift to Medium-Priority (300s):
    • Action: The load on ‘db_high_prio’ is returned to its 1x baseline level. Simultaneously, the load on a ‘db_medium_prio’ instance is increased by a factor of 30x.
    • Purpose: To test the policy’s agility and fairness. A superior policy must not only deallocate resources from the now-cold high-priority instance but also effectively reallocate them to the new, less-critical hotspot.